VISTA IOT PRELIMINARY AGENDA, 2010-01-26 14-16h Garching / 13-15h UK / 10-12h Chile 1. VISTA status and planned activities (TSz) 2. Phase II progress (MRe) 3. VISTA PPRS Summary (VDI) 4. Science Verification Data Processing Report (MAr, MRe, MPG ?, VDI) Will we be ready with a paper draft for the next Messenger? 5. Action Items review (see bellow a summary from the 2009-12-01 IOT) 6. Others 7. Date of the next meeting: end of Feb, 2010, possible mini-IOT before the start of the next "period" * * * 5: Action Items Progress report on previous action items: * AI 2009-01: Monitor execution times & survey completion status Is being followed up by MAR, MRe. 2009-07-29: There is a first prototype 2009-10-06: no news (MRe) 2009-12-01: no news STATUS: ONGOING * AI 2009-04: JEm: get in contact with F. Comeron to formalize the SADT delivery. Not yet ready to do this, though JEm chatted to Fernando about it. Documentation still missing. Discussion about the magnitude ranges that SADT takes for AG/AO stars. TSz: this is a configuration issue. Thought to be ready On September 6, 2009 2009-10-06: a report will be ready on Nov 17 (MAr) 2009-12-01: the report is not ready (MAr) STATUS: OPEN * AI 2009-06: VIV: by the end of June send the new simulation report and code to be taken over for future simulation. 2009-10-06: report writing in progress (VIv); to be transferred to the EST during Oct SciVer (TSz) 2009-12-01: MRe: the simulation delivered to OPO and Gaiteee takes over. VIv: the completeness rate is ~85%, we will need one extra year to complete the already approved surveys. TSz: we have to look again at the overheads STATUS: ONGOING * AI 2009-07: MAR/MRe: contact PIs at the mid September (once the handover is done/certain) and inform them about the time line and activities. Should wait until the Closed Cycle Cooler is repaired. 2009-10-06: done (MRe, MAr) STATUS: CLOSED * AI 2009-09: Offset sky template would be requested for the next (post SV, dry run) instrument package. Pipeline data reduction recipes are ready to work with the offset sky data sets. 2009-10-06: VIv should organize it, with help from software (i.e. Pedro Baksai), and help from JEm; by Dec 2009; starts to be changed from on hold to open (MAr) 2009-12-01: TSz: close this AI; at the moment a template with sky offsets is not needed. Backwards compatibility might be an issue. STATUS: CLOSED * AI 2009-11: TSz should find out if the constraints sets could be carried with the OBs, to the OS and the fits header of the VIRCAM files. Otherwise it is not easy to propagate these data by interfaces. 2009-07-29: TSz iterating with R. Schmutzer, but DFI (T. Bierwirth) will also need to be involved. For surveys, given the small range of programs it is expected that there are not too many different OBs, and this might need to be followed in a different/manual way in the beginning - this could imply a major change for the tools that cannot be done in the last moment. STATUS: ONGOING * AI 2009-12: TSz should check how/if the fits keywords of the raw file have the information about the concatenation and other scheduling containers. 2009-07-29: Answer is NO. But is this necessary? JEm thinks yes if processing of data takes this into account. TSz: this will need to be done together with P2PP/OT programmers and VLT software support after discussion. TSz will raise it, but cannot promise a solution. 2009-10-06: the information about the containers must be included in the keywords, CASU must be included in this work because it will have an effect on the pipeline (JEm); this is also relevant for the sky offset template (MAr); not really, because the sky offset is in the same OB as the primary target tile (JEm); in the past we used for this purpose the DPR keyworks (TSz); this info might be needed also for the ESO archive users (WHu); a good example for the archive is to identify the telluric standards, corresponding to a given data set, so the issue is relevant generally for all ESO instruments (MRe). 2009-12-01: MRe and Th. B. are looking into it, working with the BoB people. STATUS: ONGOING * AI 2009-14: GHu to check that the seeing distribution function in the scheduler for VIRCAM takes into account the finite (0.51") nominal resolution of the instrument. STATUS: ONGOING * AI 2009-16: TSz to get confirmation from A. Kaufer that the letter can be sent to SciVer and public survey PIs. 2009-10-06: done (TSz) STATUS: CLOSED * AI 2009-17: MRe should request that EVM check that SADT and p2pp were using the same instrument package. 2009-10-06: done (TSz) STATUS: CLOSED * AI 2009-18: MAR and JEm should iterate the fact sheet to be sent to the public survey PIs. 2009-10-06: done (TSz) STATUS: CLOSED * AI 2009-19: JEm should sent the latest updates for the exposure time calculator to Marina (silver reflectivity, filter curves) 2009-10-06: the transmission curves for the NB filters are still missing (MRe); OK, will send them when back to UK (JEm); next the discussion goes towards the differences between the ESO and VPO VISTA ETCs: it is important to understand the differences (MRe); the VPO ETC was not even meant to be updated anymore to avoid competition with the ESO ETC, which was meant to be the main VISTA ETC (JEm); how is the sky brightness calculated in the VPO VISTA ETC? (TSz); it is a parameter that can be modified by the user (JEm); there are known problems with the sky brightness in the IR ESO ETCs, especially in the H-band, they seem to be calculated from a wrong sky spectrum, although there is a good one available, but it is used only for the EELT ETC (TSz). 2009-12-01: done (TSz) STATUS: CLOSED * AI 2009-20: installation of a new served for faster/cheaper/better data reduction, to be available for semi-real time data reduction during the Science Verification run on 16-30.10.2009. TSz will talk to Christophe Dumas. 2009-12-01: the data had to be compressed to ensure that a copy is kept until the USD disks arrive in Garching and are checked STATUS: ONGOING * AI 2009-21: Create poor weather conditions (particularly poor seeing) back up SciVer programs. The approved SciVer programs request seeing 0.8 arcsec or better, if the SciVer was last week, it would result in almost zero data. STATUS: ONGOING * AI 2009-22: Increase the jitter box size of the sky flats to ~1/2 of the detector size. STATUS: ONGOING * * New Action Items from this meeting: * AI 2009-23: JEm should deliver the SADT by Dec 10 STATUS: NEW * AI 2009-23: MHi will contact Andine about extrapolating the 2MASS to Z-band magnitudes. STATUS: NEW * * * 6: Others N.A. * * * 7: Date of the next meeting: 26 Jan, 2010, 14 h Garching / 10 h S-go